OpenStudy (anonymous):

can you answer this question? X= log27(9). solve for X

OpenStudy (anonymous):

\[27^x=9\]solve for \(x\)

OpenStudy (anonymous):

since \(27=3^3\) and \(9=3^2\) you have \[(3^3)^x=3^2\] or \[3^{3x}=3^2\] which means \[3x=2\]